# Very Massive Stars and High N/O: A Tale of the Nitrogen-enriched Super Star Cluster in the Sunburst Arc

### v1

@@ 2.1 Abundances of pristine gas
The elemental abundances of the pristine gas are dependent on the metallicity (e.g., Berg et al., 2016, and other citations below) due to prior chemical evolution in the host galaxy. They should be modeled in this toy model as well as in simulations. Knowing that crudely $Z\propto\rm O/H$, we use oxygen abundance as a proxy for metallicity. We define $x=12+\log({\rm O/H})$, and for the solar metallicity defined in the FIRE simulations (following Wiersma et al., 2009), $x_{\odot}\approx 8.75$.
@@
We fit the data presented in Mollá et al. (2007) (galactic/extragalactic Hii regions, low metallicity galaxies; see references therein) and force the relation to match FIRE’s solar-metallicity ISM at $x=x_{\odot}$ (for compatibility with FIRE conventions), we have
@@
This relation asymptotically leads to $\log(\rm N/O)\approx-1.4$ at $Z\sim 0.2\,Z_{\odot}$. This is roughly in agreement with local observations (Vincenzo et al., 2016). Similarly, fitting data in Berg et al. (2016), we obtain
@@
This relation is not monotonic in $x$, which agrees with observations of halo stars (Fabbian et al., 2009). Our fitting relations closely match empirical relations like Nicholls et al. (2017).
Figure 2: Evolution of the relative nitrogen (top) and carbon (bottom) of the interstellar medium in the toy model (§ 2). In each panel, we show 6 tracks with various initial metallicities $Z$, and the rough estimated range of the Sunburst Arc cluster for reference (shaded regions, Pascale et al., 2023). From left to right, in each column is the model with different assumptions (closed box or leaky box, with or without VMSs of $100\,M_{\odot}$). In particular, we also show the abundance ratios in the stellar wind (i.e., without mixing with the interstellar medium; dashed lines).
2.2 Closed box vs. leaky box
Once the chemical components of the pristine gas are rescaled following the fitting relations, we are ready to study the chemical enrichment due to stellar winds. Our toy model assumes two scenarios of pristine gas evolution: either without outflow, or with feedback-driven outflow that depletes a fraction of the pristine gas from the star cluster. These scenarios correspond to classical models of “closed box” and “leaky box” respectively.
With star formation going on, the chemical components of the intra-cluster gas evolve due to enrichment from the wind ejecta, which is also metallicity-dependent. We compile evolutionary models based on $Z=0.001,0.002,0.004,0.008,0.014,0.02$ tracks of the PARSEC data. Fig. 2 shows the evolution of $\log(\rm N/O)$ and $\log(\rm C/O)$ at these metallicities, in comparison with the approximate observed ranges (gray shaded regions) of the high-pressure gas of the LyC cluster taken from Pascale et al. (2023). For the closed box model, even if VMSs are included in the stellar population, none of the evolution tracks can reach the observed N/O. For the leaky box model without the VMS population (middle column), only evolution tracks with $Z=0.014,0.02$ can marginally reach the lower bound of the observed range at late times, but their $\log(\rm C/O)$ is well above the observed range. Even if we consider the extreme case that the leftover gas in the cluster is purely ejecta-dominated (dashed lines), the C/O ratio is well above the observation.
@@
The carbon abundance is another constraint to be satisfied. Based on it, Tapia et al. (2024) argues that WR stars should be excluded from the model to avoid over-production of carbon. However, the PARSEC model here suggests that carbon enrichment can be modest to match the observation for certain metallicities (i.e., $Z=0.002$ and $Z=0.004$ here). Interestingly, these metallicities are roughly equal to $0.2\,Z_{\odot}$, the metallicity of the LyC cluster (Pascale et al., 2023). This further supports the “leaky box + VMS model” as a possible explanation for this specific system.
2.3 Summary and caveats
@@
This toy model cannot capture all the complexity of chemical enrichment during star cluster formation, as it omits several key physical processes. First, it lacks a realistic star formation history, which would influence the timing and intensity of enrichment. Second, it neglects gas migration relative to the cluster, whereas the true gas distribution is often far more complex than the assumed homogeneous configuration (Pascale et al., 2023). Third, the model does not account for the inhomogeneous nature of chemical feedback, wherein gas parcels surrounding individual stars can be enriched to different degrees. Finally, the adopted outflow prescription ignores supernova explosions and their associated feedback, which can inject substantial energy after the starburst phase and rapidly expel gas from the cluster. Lifting these limitations will require radiation-hydrodynamic simulations of star formation, which can capture the spatial and temporal complexities of gas dynamics and feedback, thereby providing a more realistic picture of chemical enrichment during cluster assembly. We present such simulations in the following sections.
3 Simulations
@@
We run magnetohydrodynamical (MHD) simulations of star formation with the code GIZMO (Hopkins, 2015) in its meshless finite-mass (MFM) mode, which solves the MHD equations in Hopkins & Raives (2016). Star formation is modeled with FIRE-3 (Hopkins et al., 2014, 2018a, 2023), incorporating additional physics like $N$-body gravity, radiation transfer, photo-heating, photo-ionization, photon momentum, mechanical outflows, etc.
@@
The underlying IMF of FIRE-3 does not cover the VMS regime, so we implemented modifications to the standard FIRE physics to model the effects of VMSs (as in Shi & Murray, 2025). We first quickly recap this framework.
3.1 Hybrid FIRE and VMS method
@@
We use a modified version of FIRE star formation physics, which, in addition to converting gas cells into “stars” representing a single stellar population (SSP), also converts a fraction of the mass into individually resolved stars and models their evolution in greater detail. As such, there are two kinds of “stars” in the simulation: FIRE SSP particles representing relatively faint stars below some cutoff $m_{\rm cut}$, and individual massive stars above the cutoff. This is described in Shi & Murray (2025).
@@
This implementation has two major aspects. The first aspect is the inclusion of the VMS population. In FIRE-2 (Hopkins et al., 2018a) and FIRE-3 (Hopkins et al., 2023) models, the stellar evolution with population synthesis models and the IMF’s high-mass end cutoff is $100\,M_{\odot}$, so the effect of more massive stars is not included. With the hybrid method, we presume an IMF with a higher mass cutoff (e.g., $300\,M_{\odot}$) and distribute the newly-formed stellar mass into the traditional FIRE-like SSP particles and individual VMS particles above some cutoff $m_{\rm cut}$. These individual VMS particles will be modeled in more detail.
@@
Another aspect is that the method can simulate massive GMCs of $\sim 10^{8}\,M_{\odot}$ more cost-efficiently yet tracks individual VMSs. Sophisticated star formation simulation modules like STARFORGE (Grudić et al., 2021; Guszejnov et al., 2021) model the formation of individual stars self-consistently but typically require very high mass resolutions ($\sim 0.01\,M_{\odot}$) and can be extremely computationally expensive to run GMCs of $\sim 10^{8}\,M_{\odot}$. On the other hand, FIRE simulations are typically much faster due to the lower resolution ($\sim 100-1000\,M_{\odot}$) but lack the modeling of star-by-star physics. The hybrid method can model the evolution of individual VMSs and their radiative and mechanical feedback in the same way as FIRE SSP particles (Hopkins et al., 2018b, 2020). In practical simulations, we find that the additional cost to simulate VMSs compared to purely FIRE simulations is not significant unless an enormous number of VMS particles are added to the simulations, which causes memory load-balance issues.
@@
In summary, the hybrid method is suitable for studying star formation problems where VMSs play an important role, e.g., the formation and growth of massive remnant BHs of VMSs (Shi & Murray, 2025). As proved in § 2 and Fig. 2, VMSs dominate the chemical feedback by stellar winds; it is then proper to apply the method to simulate this process.
@@
Figure 3: Mass evolution of VMSs compiled from the PARSEC results (Costa et al., 2025). Here we present the evolution tracks (solid) for stars with ZAMS mass between $100\,M_{\odot}$ and $300\,M_{\odot}$ and $Z=0.004,0.014$. We also show fittings of these evolution tracks before (following Eq. 5; dashed) and after (linear relation; dotted) the “kink.”
3.2 Evolution of Very Massive Stars
@@
We update the VMS evolution model originally presented in Shi & Murray (2025) with the PARSEC data set. In particular, we modify the mass evolution recipe from Vink (2018) to a new model directly compiled from the PARSEC data. We find the mass evolution of stars in PARSEC typically has two phases, an early hydrogen-burning phase with a lower mass loss rate, and the succeeding phase with a higher mass loss rate, yielding a “kink” in the mass-time plot (Fig. 3). The surface chemical abundances $X_{i}$ also change significantly at $t_{\rm kink}$. We find the mass loss rate before $t_{\rm kink}$ is well fitted with
@@
The PARSEC evolution tracks are in agreement with other studies of VMS evolution, that they can lose significant mass in winds before the end of their lifetime Vink (2018, 2023, 2024). Following the improved mass-loss recipe (Vink et al., 2011), Vink (2018) found that the terminal wind velocity of VMSs is $\lesssim 500\,\rm km/s$, significantly slower than extrapolations from the $100\,M_{\odot}$ massive stars (Leitherer et al., 1992; Lamers et al., 1995). Unlike the mass loss rate, the wind velocity is not tightly correlated with the stellar mass. In the simulation, we model the wind velocity $v_{\rm wind}$ with a fitted linear relation based on Vink (2018):
@@
Compared with predictions in Vink (2018), the scatter is $\lesssim 50\,\rm km/s$ in the mass regime of our interest (i.e., $100-300\,M_{\odot}$).
@@
In the supernova stage after the main sequence, we distribute the mass in supernova winds into 11 elements following an interpolation table compiled from PARSEC (as a function of $m_{0}$ and $Z$). We also note that not all VMSs end up with a supernova explosion  (Heger et al., 2003; Costa et al., 2025) and this leads to some modifications to the feedback from FIRE SSP particles (§ 3.3).
3.3 Changes to the FIRE-3 physics
Figure 4: The SNe chemical yield of massive stars and very massive stars compiled from the PARSEC data set, for stars with initial metallicity $Z=0.004$. Top panel - the SNe yield as a function of the zero-age main-sequence (ZAMS) stellar mass for selected elements (C, N, O, Mg, Fe); note that the fate of a star of a given metallicity is dependent on its ZAMS mass (color shaded regions), including core-collapse SN (CCSN), failed SN (FSN), pulsational pair-instability SN (PPISN), and direct-collapse black hole (DCBH). Bottom panel - the SNe yield as a function of stellar age.
We make some subtle changes to the standard FIRE-3 physics (Hopkins et al., 2023) in two aspects.
@@ 3.3 Changes to the FIRE-3 physics
Moreover, we modify the supernovae (SNe) feedback model of FIRE-3, which includes the impact from both core-collapse SNe and SNe Ia of each FIRE SSP particle (Hopkins et al., 2023). The early stage of the SNe feedback in FIRE-3 is dominated by core-collapse SNe once the age of the stellar population is older than $3.7\,\rm Myr$. However, from more detailed stellar evolution simulations with PARSEC in Costa et al. (2025), massive stars of $\sim 30$$75\,M_{\odot}$ (depending on metallicity) may only terminate as “failed SNe” that inject little mass and energy into the interstellar medium due to their strong self-gravity, as indicated by theory (Heger et al., 2003) and observations (e.g., N6946-BH1, Gerke et al., 2015; Beasor et al., 2024). Instead, (“successful”) core-collapse SNe can only appear after $t_{\rm CC}\sim 7$$7.8\,\rm Myr$ (depending on metallicity), corresponding to the lifetime of $m\sim 30\,M_{\odot}$ stars. We find $t_{\rm CC}$ can be well fitted with
@@
Fiducial simulations (§ 4.1) $3\times 10^{7}$ 60, 70, 80 0.15, 0.2, 0.25 FIRE+VMS
GMC surface density tests (§ 4.2) $10^{7},10^{8}$ 60, 70, 80 0.2 FIRE+VMS
FIRE-only simulations (§ 4.3) $3\times 10^{7}$ 60, 70, 80 0.2 FIRE (unaltered, or with FSN)
VMS feedback tests (§ 4.4) $3\times 10^{7}$ 60, 70, 80 0.2 FIRE+VMS (disabling feedback)
@@
3.4 Simulation suite
Finally, we describe other details of the simulation runs performed in this study.

### v2

@@ II.1 Abundances of pristine gas
The elemental abundances of the pristine gas are dependent on the metallicity (e.g., BergSkillmanHenry_2016ApJ...827..126B, and other citations below) due to prior chemical evolution in the host galaxy. They should be modeled in this toy model as well as in simulations. Knowing that crudely $Z\propto\rm O/H$, we use oxygen abundance as a proxy for metallicity. We define $x=12+\log({\rm O/H})$, and for the solar metallicity defined in the FIRE simulations (following WiersmaSchayeSmith_2009MNRAS.393...99W), $x_{\odot}\approx 8.75$.
Figure 2: textcolorblackFitting relations of log(N/O) (Eq. 3; top) and log(C/O) (Eq. 4; bottom), which match the solar metallicity used in FIRE (WiersmaSchayeSmith_2009MNRAS.393...99W; HopkinsWetzelWheeler_2023MNRAS.519.3154H). We also present data points for log(N/O) (MollaVilchezDiaz_2007astro.ph..1691M, and references therein) and log(C/O) (BergSkillmanHenry_2016ApJ...827..126B, and references therein), where the low-redshift LyC leaking galaxies (red dots) are compiled from IzotovSchaererWorseck_2023MNRAS.522.1228I.
@@
We fit the data presented in MollaVilchezDiaz_2007astro.ph..1691M (galactic/extragalactic Hii regions, low metallicity galaxies; see references therein) and force the relation to match FIRE’s solar-metallicity ISM at $x=x_{\odot}$ (for compatibility with FIRE conventions), we have
@@
This relation asymptotically leads to $\log(\rm N/O)\approx-1.4$ at $Z\sim 0.2\,Z_{\odot}$. This is roughly in agreement with local observations (VincenzoBelfioreMaiolino_2016MNRAS.458.3466V). Similarly, fitting data in BergSkillmanHenry_2016ApJ...827..126B, we obtain
@@
This relation is not monotonic in $x$, which agrees with observations of halo stars (FabbianNissenAsplund_2009A&A...500.1143F).
textcolorblackFig. 2 presents the fitting relations (Eqs. 3 and 4) along with data from MollaVilchezDiaz_2007astro.ph..1691M and BergSkillmanHenry_2016ApJ...827..126B. Given the large scattering of the raw data, our fitting relations approximately match empirical relations in NichollsSutherlandDopita_2017MNRAS.466.4403N. We also plot the log(N/O) and log(C/O) of the Sunburst Arc LyC cluster (PascaleDaiMcKee_2023ApJ...957...77P), which suggests anomalously elevated N/O, while C/O is not significantly elevated.
Figure 3: Evolution of the relative nitrogen (top) and carbon (bottom) of the interstellar medium in the toy model (§ II). In each panel, we show 6 tracks with various initial metallicities $Z$, and the rough estimated range of the Sunburst Arc cluster for reference (shaded regions, PascaleDaiMcKee_2023ApJ...957...77P). From left to right, in each column is the model with different assumptions (closed box or leaky box, with or without VMSs of $100\,M_{\odot}$). In particular, we also show the abundance ratios in the stellar wind (i.e., without mixing with the interstellar medium; dashed lines).
II.2 Closed box vs. leaky box
Our toy model considers two scenarios for the pristine gas: one without outflow and one with feedback-driven outflow that removes a fraction of the gas from the cluster. These correspond to the classical “closed-box” and “leaky-box” models, respectively.
The chemical components of the polluted gas evolve due to enrichment from the wind ejecta. Fig. 3 shows the evolution of $\log(\rm N/O)$ and $\log(\rm C/O)$ at metallicities $Z=0.001,0.002,0.004,0.008,0.014,0.02$, in comparison with the approximate observed ranges (gray shaded regions) of the high-pressure gas of the LyC cluster taken from PascaleDaiMcKee_2023ApJ...957...77P. For the closed box model, even if VMSs are included in the stellar population, none of the evolution tracks can reach the observed N/O. For the leaky box model without the VMS population (middle column), only evolution tracks with $Z=0.014,0.02$ can marginally reach the lower bound of the observed range at late times, but their $\log(\rm C/O)$ is well above the observed range. Even if we consider the extreme case that the leftover gas in the cluster is purely ejecta-dominated (dashed lines), the C/O ratio is well above the observation.
@@
The carbon abundance is another constraint to be satisfied. Based on it, TapiaBekkiGroves_2024MNRAS.534.2086T argues that WR stars should be excluded from the model to avoid over-production of carbon. However, the PARSEC model here suggests that carbon enrichment can be modest to match the observation for certain metallicities (i.e., $Z=0.002$ and $Z=0.004$ here). Interestingly, these are roughly $0.2\,Z_{\odot}$, the metallicity of the LyC cluster (PascaleDaiMcKee_2023ApJ...957...77P). This further supports the “leaky box + VMS model” as a possible explanation for this specific system.
textcolorblackIt is worth mentioning the two low-metallicity tracks ($Z=0.001,10^{-4}$) also overproduce carbon after $\sim 3\,\rm Myr$. This is primarily due to the “dredge-up events” happening to a fraction of massive stars near the end of their time, where the convective envelope penetrates to the helium-burning core and brings the carbon to the surface (Appendix C of CostaShepherdBressan_2025A&A...694A.193C). This could affect the chemical feedback in systems with even lower metallicity than discussed here.
II.3 Summary and caveats
@@
This toy model omits several key physical processes. First, it lacks a realistic star formation history, which affects the timing and strength of enrichment. Second, it neglects gas migration relative to the cluster, although real gas distributions are far more complex than the assumed homogeneous one (PascaleDaiMcKee_2023ApJ...957...77P). Third, it ignores the inhomogeneous nature of chemical feedback, where gas around individual stars can be enriched to varying degrees. Finally, the outflow prescription excludes supernova explosions and their feedback, which can inject significant energy after the starburst and rapidly deplete cluster gas. Addressing these limitations requires radiation-hydrodynamic simulations that capture the spatial and temporal complexities of feedback and gas dynamics, offering a more realistic view of chemical enrichment during cluster assembly. We present such simulations below.
III Simulations
@@
We run magnetohydrodynamical (MHD) simulations of star formation with the code GIZMO (Hopkins_2015MNRAS.450...53H) in its meshless finite-mass (MFM) mode, which solves the MHD equations in HopkinsRaives_2016MNRAS.455...51H. Star formation is modeled with FIRE-3 (HopkinsKeresOnorbe_2014MNRAS.445..581H; HopkinsWetzelKeres_2018MNRAS.480..800H; HopkinsWetzelWheeler_2023MNRAS.519.3154H), incorporating additional physics like $N$-body gravity, radiation transfer, photo-heating, photo-ionization, photon momentum, mechanical outflows, etc.
@@
The underlying IMF of FIRE-3 does not cover the VMS regime, so we implemented modifications to the standard FIRE physics to model the effects of VMSs (as in Shi_inprep). We first quickly recap this framework.
III.1 Hybrid FIRE and VMS method
@@
We employ a modified version of the FIRE star formation model that, in addition to converting gas cells into “stars” representing single stellar populations (SSPs), also forms individually resolved massive stars whose evolution is followed in detail. Thus, the simulation contains two types of “stars”: FIRE SSP particles representing less-massive stars below a cutoff mass $m_{\rm cut}$, and individual massive stars above it. A detailed description is given in Shi_inprep.
@@
This implementation has two main features. The first is the inclusion of a very massive star (VMS) population. In the FIRE-2 (HopkinsWetzelKeres_2018MNRAS.480..800H) and FIRE-3 (HopkinsWetzelWheeler_2023MNRAS.519.3154H) models, stellar evolution and population synthesis adopt an IMF cutoff of $100\,M_{\odot}$, excluding more massive stars. Our hybrid method extends this to higher cutoffs (e.g., $300\,M_{\odot}$), distributing newly formed stellar mass between traditional FIRE-like SSP particles and individual VMS particles above a chosen cutoff $m_{\rm cut}$, which are evolved in detail.
@@
The second feature enables simulations of massive GMCs ($\sim 10^{8}\,M_{\odot}$) with improved efficiency while still resolving individual VMSs. Modules such as STARFORGE (GrudicGuszejnovHopkins_2021MNRAS.506.2199G; GuszejnovGrudicHopkins_2021MNRAS.502.3646G) model individual stars self-consistently but require extremely high resolution ($\sim 0.01\,M_{\odot}$) and are computationally costly for such large GMCs. By contrast, FIRE simulations are faster at lower resolution ($\sim 100$$1000\,M_{\odot}$) but lack star-by-star physics. The hybrid method bridges this gap, following the evolution and feedback of individual VMSs analogously to FIRE SSP particles (HopkinsWetzelKeres_2018MNRAS.477.1578H; HopkinsGrudicWetzel_2020MNRAS.491.3702H). In practical simulations, we find that the additional cost to simulate VMSs is insignificant unless an enormous number of VMS particles are created, which causes memory load-balance issues.
@@
In summary, the hybrid method is suitable for studying star formation problems where VMSs play an important role, e.g., the formation and growth of massive remnant BHs of VMSs (Shi_inprep). As proved in § II and Fig. 3, VMSs dominate the chemical feedback by stellar winds; it is then proper to apply the method to simulate this process.
@@
Figure 4: Mass evolution of VMSs compiled from the PARSEC results (CostaShepherdBressan_2025A&A...694A.193C). Here we present the evolution tracks (solid) for stars with ZAMS mass between $100\,M_{\odot}$ and $300\,M_{\odot}$ and $Z=0.004,0.014$. We also show fittings of these evolution tracks before (following Eq. 5; dashed) and after (linear relation; dotted) the “kink.”
III.2 Evolution of VMSs
@@
We update the VMS mass-evolution prescription of Shi_inprep using a new model compiled directly from the PARSEC data. textcolorblackOwing to the bistability jump (VinkdeKoterLamers_1999A&A...350..181V), the evolution first proceeds through an early hydrogen-burning phase with a low mass-loss rate, followed by a phase with a higher rate, producing a distinct “kink” in the mass–time relation (Fig. 4). The surface abundances $X_{i}$ are also different in these phases. The mass-loss rate before $t_{\rm kink}$ is well fitted with
@@
The PARSEC evolution tracks are in agreement with other studies of VMS evolution, that they can lose significant mass in winds before the end of their lifetime Vink_2018A&A...615A.119V; Vink_2023A&A...679L...9V; Vink_2024arXiv241018980V. Following the improved mass-loss recipe (VinkMuijresAnthonisse_2011A&A...531A.132V), Vink_2018A&A...615A.119V found that the terminal wind velocity of VMSs is $\lesssim 500\,\rm km/s$, significantly slower than extrapolations from the $100\,M_{\odot}$ massive stars (LeithererRobertDrissen_1992ApJ...401..596L; LamersSnowLindholm_1995ApJ...455..269L). Unlike the mass loss rate, the wind velocity is not tightly correlated with the stellar mass. In the simulation, we model the wind velocity $v_{\rm wind}$ with a fitted linear relation based on Vink_2018A&A...615A.119V:
@@
Compared with predictions in Vink_2018A&A...615A.119V, the scatter is $\lesssim 50\,\rm km/s$ in the mass regime of our interest (i.e., $100-300\,M_{\odot}$).
@@
In the supernova stage after the main sequence, we distribute the mass in supernova winds into 11 elements following an interpolation table compiled from PARSEC (as a function of $m_{0}$ and $Z$). We also note that not all VMSs end up with a supernova explosion  (HegerFryerWoosley_2003ApJ...591..288H; CostaShepherdBressan_2025A&A...694A.193C) and this leads to some modifications to the feedback from FIRE SSP particles (§ III.3).
III.3 Changes to the FIRE-3 physics
Figure 5: The SNe chemical yield of massive stars and VMSs compiled from the PARSEC data set, for stars with initial metallicity $Z=0.004$. Top panel - the SNe yield as a function of the zero-age main-sequence (ZAMS) stellar mass for selected elements (C, N, O, Mg, Fe); note that the fate of a star of a given metallicity is dependent on its ZAMS mass (color shaded regions), including core-collapse SN (CCSN), failed SN (FSN), pulsational pair-instability SN (PPISN), pair-instability SN (PISN), and direct-collapse black hole (DCBH). Bottom panel - the SNe yield as a function of stellar age.
@@ III.3 Changes to the FIRE-3 physics
Moreover, we modify the supernovae (SNe) feedback model of FIRE-3, which includes the impact from both core-collapse SNe and SNe Ia of each FIRE SSP particle (HopkinsWetzelWheeler_2023MNRAS.519.3154H). The early stage of the SNe feedback in FIRE-3 is dominated by core-collapse SNe once the age of the stellar population is older than $3.7\,\rm Myr$. However, from more detailed stellar evolution simulations with PARSEC in CostaShepherdBressan_2025A&A...694A.193C, massive stars of $\sim 30$$75\,M_{\odot}$ (depending on metallicity) may only terminate as “failed SNe” that inject little mass and energy into the interstellar medium due to their strong self-gravity, as indicated by theory (HegerFryerWoosley_2003ApJ...591..288H) and observations (e.g., N6946-BH1, GerkeKochanekStanek_2015MNRAS.450.3289G; BeasorHosseinzadehSmith_2024ApJ...964..171B). Instead, (“successful”) core-collapse SNe can only appear after $t_{\rm CC}\sim 7$$7.8\,\rm Myr$ (depending on metallicity), corresponding to the lifetime of $m\sim 30\,M_{\odot}$ stars. We find $t_{\rm CC}$ can be well fitted with
@@
Fiducial simulations (§ IV.1) $3\times 10^{7}$ 60, 70, 80 0.15, 0.2, 0.25 FIRE+VMS
GMC surface density tests (§ IV.2) $10^{7},10^{8}$ 60, 70, 80 0.2 FIRE+VMS
FIRE-only simulations (§ IV.3) $3\times 10^{7}$ 60, 70, 80 0.2 FIRE (unaltered, or with FSN)
VMS feedback tests (§ IV.4) $3\times 10^{7}$ 60, 70, 80 0.2 FIRE+VMS (disabling feedback)
@@
III.4 Simulation suite